# On Sensitivity of Learning with Limited Labelled Data to the Effects of Randomness: Impact of Interactions and Systematic Choices

### v1

@@ Abstract
While learning with limited labelled data can improve performance when the labels are lacking, it is also sensitive to the effects of uncontrolled randomness introduced by so-called randomness factors (e.g., varying order of data). We propose a method to systematically investigate the effects of randomness factors while taking the interactions between them into consideration. To measure the true effects of an individual randomness factor, our method mitigates the effects of other factors and observes how the performance varies across multiple runs. Applying our method to multiple randomness factors across in-context learning and fine-tuning approaches on 7 representative text classification tasks and meta-learning on 3 tasks, we show that: 1) disregarding interactions between randomness factors in existing works caused inconsistent findings due to incorrect attribution of the effects of randomness factors, such as disproving the consistent sensitivity of in-context learning to sample order even with random sample selection; and 2) besides mutual interactions, the effects of randomness factors, especially sample order, are also dependent on more systematic choices unexplored in existing works, such as number of classes, samples per class or choice of prompt format.
@@
Branislav ${}^{\spadesuit}$${}^{\dagger}$${}^{\ddagger}$, Ivan ${}^{\dagger}$, Maria ${}^{\dagger}$${}^{\ddagger}$$^{\spadesuit}$ Faculty of Information Technology, Brno University of Technology, Brno, Czechia ${}^{\dagger}$ Kempelen Institute of Intelligent Technologies, Bratislava, Slovakia ${}^{\ddagger}$ Slovak.AI, Bratislava, Slovakia {branislav.pecher, ivan.srba, maria.bielikova}@kinit.sk
@@ 1 Introduction
Learning with limited labelled data is an umbrella term for approaches designed to work when enough labels are lacking. This includes approaches such as in-context learning, where a language model is prompted using a concatenation of task instruction, a few examples and a test sample, and its output is mapped to a label Liu et al. (2023), fine-tuning, where a pre-trained language model is trained for a few steps on a supervised downstream task Min et al. (2021), or meta-learning, where the model learns how to effectively adapt to multiple tasks at the same time Hospedales et al. (2021).
Although such approaches can effectively deal with limited labels, they were observed to be notably sensitive to the effects of uncontrolled randomness, which negatively affects their stability. Under stability, we understand a property of a learning algorithm or a trained model that indicates what influence the small scale and random changes in the input data and parameters have on its outputs Liu et al. (2017). The randomness in the training process can have massive impact, leading to large deviation in the performance over multiple training runs. In-context learning was found to be sensitive to the order of samples, where changing this order can lead from state-of-the-art predictions to random guessing Lu et al. (2022); Zhao et al. (2021b). Repeating fine-tuning multiple times can also result in smaller language models outperforming their larger counterparts Dodge et al. (2020), or based on a choice of adaptation samples in meta-learning, the performance can vary between 4% or 95% Agarwal et al. (2021). The randomness is introduced by the randomness factors, which represent the non-deterministic decisions in the training process, such as order of samples, the model initialisation or sample choice Pham et al. (2021); Gundersen et al. (2022); Pecher et al. (2023).
@@
If the randomness is not properly addressed, it can have non-negligible negative consequences, even with enough labelled samples Reimers and Gurevych (2017); Dodge et al. (2020); McCoy et al. (2020). It can prohibit objective comparison with the state-of-the-art baselines, with a method being incorrectly denoted as state-of-the-art only based on more favourable chance Reimers and Gurevych (2017). The uncontrolled randomness can also intentionally (but, unfortunately, also intentionally by cherry-picking) create an imaginary perception of research progress. It was also identified as a major obstacle to reproducibility Albertoni et al. (2023).
@@
A lot of focus is dedicated to investigating and mitigating the effects of randomness on the stability of learning with limited labelled data Mosbach et al. (2021); Zhao et al. (2021a); Pecher et al. (2023), especially for in-context learning Lu et al. (2022); Zhao et al. (2021b); Chang and Jia (2023); Li and Qiu (2023); Gao et al. (2021); Köksal et al. (2023), while often being limited in its extent (focusing on a small fragment of factors or approaches) and at times leading to contradictory results. For example, in-context learning was long believed to be consistently sensitive to the order of the randomly selected samples Lu et al. (2022); Zhao et al. (2021b). However, it was later observed that this sensitivity disappears when a more sophisticated sample selection strategy was used Zhang et al. (2022); Chang and Jia (2023); Li and Qiu (2023). Therefore, the current research does not clearly answer which of these two interconnected randomness factors (sample order and sample selection) is an actual source of performance deviations.
@@
We argue that such observed inconsistencies are caused by disregarding the interactions between randomness factors, which leads to incorrectly attributing the performance deviations to different randomness factors. Such interactions are so far partially or even completely overlooked in the existing works, contributing to the misleading findings. In addition, we hypothesise that the sensitivity of in-context learning is not only affected by the interactions with other randomness factors but also by other systematic choices, which are not thoroughly controlled and explored in the existing works, such as the number of classes, shots per class and prompt format.
@@
Our main contributions are as follows [^1] To support replicability and extension of our results, we openly publish the source code of our proposed method at https://anonymous.4open.science/r/investigation-with-interactions-F6D4:
@@
  • We propose a novel method for investigation of randomness factors’ effects that, in contrast to the existing works, explicitly address interactions between them by mitigating the effects of other non-investigated factors. In addition, it measures the relative importance of the effects of randomness factors by calculating what fraction of the overall deviation in the model’s performance (estimated by a golden model) the investigated factor contributes in comparison to all other factors.
@@
  • We further analyse how more systematic choices influence the importance of the randomness factors. We find the following key insights: 1) predicting a higher number of classes leads to increased importance of sample order for in-context learning and reduced importance of sample order and model initialisation for fine-tuning approaches; 2) increasing the number of in-context samples reduces the importance of sample selection while having no consistent effect on the importance of sample order; and 3) the choice of prompt format has a significant impact on the importance of different factors, with larger models showing lower sensitivity to this choice.
@@ 2 Related Work
The effects of randomness factors are investigated either by repeating the training multiple times without explicitly changing the training setup, relying on the non-determinism of the training process McCoy et al. (2020); Zhou et al. (2020); Dodge et al. (2020), or by explicitly changing specific aspects of training, such as changing what data is used Bouthillier et al. (2019); Pham et al. (2021); Agarwal et al. (2021); Zhao et al. (2021a). As such investigation may be affected by the interactions with other factors, another possibility is to perform the investigation by fixing all the other factors to a specific state, either chosen randomly Boquet et al. (2019); Pham et al. (2021); Zhao et al. (2021a) or as a result of a mitigation strategy Li and Qiu (2023); Chang and Jia (2023). Another investigation strategy is to vary all the investigated factors at the same time and then decouple their effects in evaluation Dodge et al. (2020); Bouthillier et al. (2021); Sellam et al. (2022), where accounting for as many factors as possible leads to more reliable results, but with a trade-off of a significant increase in computation costs Bouthillier et al. (2021).
@@
Most of the focus on investigating and mitigating effects of randomness is on in-context learning, which was found to be especially sensitive to the choice of samples Liu et al. (2022); Zhang et al. (2022); Chang and Jia (2023); Li and Qiu (2023); Köksal et al. (2023) and their order Lu et al. (2022); Zhao et al. (2021b); Nguyen and Wong (2023). However, it was observed that the sensitivity to sample order disappears when using a more sophisticated sample selection strategy instead of random selection Zhang et al. (2022); Chang and Jia (2023); Li and Qiu (2023), hinting at interactions between these factors that may lead to inconsistent results. In addition, the performance of in-context learning was found to be sensitive to more systematic choices as well Weber et al. (2023), such as the format of the prompt Sclar et al. (2023); Voronov et al. (2024) or number of shots Liu et al. (2022); Mavromatis et al. (2023). However, the impact of these systematic choices on the effects of randomness factors is not thoroughly investigated. Besides in-context learning, the language model fine-tuning approaches were identified to be sensitive to the random seeds, which influence the model initialisation and order of samples in training Dodge et al. (2020); McCoy et al. (2020); Zhou et al. (2020); Mosbach et al. (2021); Zhao et al. (2021a); Zhong et al. (2021); Sellam et al. (2022). Finally, meta-learning, investigated only by a small fraction of works, was found to be sensitive to the choice of samples for adaptation and training Agarwal et al. (2021); Setlur et al. (2021); Cioba et al. (2022) or how the samples are split into tasks Ye et al. (2021). The overall machine learning approaches were found to be sensitive to other factors as well, such as the impact of framework and hardware implementation Boquet et al. (2019); Pham et al. (2021), which was found to be lower than other factors, or the data split Bouthillier et al. (2019, 2021), with no consistent findings regarding its importance. These factors are therefore mostly considered in specific cases such as in benchmarking studies Ye et al. (2021); Bragg et al. (2021).
@@
In majority of the cases, the effects of randomness are evaluated by simply aggregating results over multiple runs and reporting a single value (e.g., mean, standard deviation or a number of failed runs Bouthillier et al. (2019); McCoy et al. (2020); Setlur et al. (2021); Ye et al. (2021); Cioba et al. (2022)), or multiple aggregated values (such as the worst, the best and the average performance, or the difference between worst and best run Agarwal et al. (2021); Zhang et al. (2022)). In specific cases, statistical approaches are used, such as grouping runs and aggregating on group level Dodge et al. (2020), decoupling interactions Boquet et al. (2019) or estimating distribution from lower number of training runs Bouthillier et al. (2021); Sellam et al. (2022). However, almost no studies analyse the importance of the effects, such as what fraction of the overall variance the specific factor contributes.
@@ 3 Investigation of Randomness while Taking Interactions into Consideration
We propose a new method for investigating the effects of any randomness factor that takes the interactions between the effects of other factors into consideration. At the same time, the method is designed to measure the importance of the found effects. The full formalisation of the method and further details are included in Appendix B, with the steps compiled in Algorithm 1.
At its core, the investigation is done by observing how the performance changes across different states (or configurations, such as the different permutations of samples in Data Order randomness factor) the investigated randomness factors can appear in (similarly to previous works). In a single investigation run, the training and evaluation of a model is done multiple ($N$) times, each time with a different configuration of the selected randomness factor (e.g., a different split of data, choice of samples or their order), while keeping the configuration of all the other non-investigated randomness factors fixed (lines 5-7 in the Algorithm 1). The standard deviation of a single investigation run (line 9 in the Algorithm 1) represents the effects of the randomness factor that are still affected by the interactions (called partial standard deviation).
To mitigate the effects of interactions, the investigation run is repeated multiple ($M$) times with different fixed configurations of the non-investigated randomness factors (lines 4-10 in the Algorithm 1). Each such repeat is called mitigation run. After performing enough mitigation runs (i.e., searching through enough configurations of non-investigated randomness factors), the partial standard deviations are averaged to produce the contributed standard deviation ($c\_std$; line 11 in the Algorithm 1), which represents the final adjusted effects of investigated randomness factor (i.e., it represents the deviation the investigated randomness factor contributes to the overall deviation in results).
To assess the importance of the factor, the contributed standard deviation is compared with two additional values: 1) mitigated standard deviation ($m\_std$); and 2) golden model standard deviation ($gm\_std$). The mitigated standard deviation represents the joint effects of all the non-investigated randomness factors and is calculated as the standard deviation across the mitigation runs (i.e., the deviation contributed by the non-investigated factors; lines 8 and 12 in the Algorithm 1). The golden model standard deviation represents an objective estimate of the deviation in the model performance and is calculated as the deviation across multiple ($L$) mitigation runs without any investigated randomness factors (e.g., repeating evaluation while changing configurations for all factors; lines 13-17 in the Algorithm 1).
The importance score of the factor is defined as the portion of the golden model standard deviation the investigated factors contribute over the non-investigated ones ($importance=(c\_std-m\_std)/gm\_std$). As the number of mitigated factors is significantly larger than the number of investigated factors, any randomness factor with an importance value over 00 (i.e., representing the state where the investigated factor contributes the same amount of deviation as all the non-investigated factors combined) is considered significantly important (lines 18-21 in the Algorithm 1).
The number of investigation runs ($N$) and the mitigation runs ($M$) provide a trade-off between the feasibility (or computation costs) of the investigation and the precision (or reliability) of the investigation results. As a result of an ablation study of this trade-off, which is included in Appendix C, the number of the runs should adhere to the following conditions to achieve a reliable estimate of the effects of randomness factors and their importance:
  1. 1.
    $N,M\gg 1$ – to guarantee the reliability of the results, $N$ and $M$ should be chosen to cover a large enough number of factor configurations.
  2. 2.
    $M\geq N$ – as the size $M$ represents a larger state space (dependent on the number of randomness factors), it should be at least equal to the number of investigation runs to mitigate the interactions, with higher difference leading to more precise results.
  3. 3.
    $L=N*M$ – to keep the importance of effects unbiased, the number of runs in the golden model should be the same as the number of the overall runs in the investigation.
4 Experiments
Datasets.
The experiments are conducted on 7 text classification datasets composed of different tasks with different number of classes. We focus on 3 binary classification datasets from the GLUE benchmark Wang et al. (2018): SST2 Dankers and Titov (2022) for sentiment classification, CoLA Warstadt et al. (2019) for determining the grammatical acceptability of a sentence, and MRPC Dolan and Brockett (2005) for determining the semantic equivalence relationship between two sentences. In addition, we use 4 multi-class text datasets: AG News Zhang et al. (2015) for news classification, TREC Voorhees and Tice (2000) for question classification, DB-Pedia Lehmann et al. (2015) for topic classification and SNIPS Coucke et al. (2018) for intent classification.
Approaches.
The main focus of the investigation is on the in-context learning using the Flan-T5 Chung et al. (2022) base model, LLaMA-2 Touvron et al. (2023) 13B instruction optimised model, Mistral-7B Jiang et al. (2023) and Zephyr-7B Tunstall et al. (2023). In addition, we also focus on fine-tuning, using the BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019) base models. Finally, we also investigate the meta-learning approaches, specifically MAML Finn et al. (2017), Reptile Nichol et al. (2018) and the Prototypical Networks Snell et al. (2017), but only on binary classification datasets.
Randomness Factors.
Method Setup.
For each randomness factor, the number of the investigation runs ($N$) is set to 10, and the number of mitigation runs is set to 100 for fine-tuning, meta-learning and Flan-T5 model and 20 for the remaining models. These numbers, selected based on an ablation study (included in Appendix 3), provide a sufficient coverage of the configurations’ state space and produce a good balance between the reliability of the results and computation costs. To keep the comparisons between the golden model and the investigated factors as unbiased as possible, we use the same overall number of runs (1 000 for fine-tuning, meta-learning and Flan-T5, 200 for the rest).
Experimental Setup.
Flan-T5 Random Fixed Interactions
Golden Model 2.244 2.244 2.244
Label Select. (*) 2.517 (*) 2.594 (*) 2.128
Data Split (*) 2.362 (*) 2.480 (*) 2.167
Data Order (*) 2.131 (*) 3.014 0.869
Sample Choice (*) 2.370 (*) 3.191 (*) 2.123
Zephyr-7B Random Fixed Interactions
Golden Model 1.043 1.043 1.043
Label Select. (*) 1.122 (*) 1.004 (*) 0.863
Data Split (*) 1.185 0.402 (*) 0.664
Data Order (*) 1.138 (*) 0.957 0.456
Sample Choice (*) 1.052 0.406 (*) 0.744
Table 1: Comparison of different investigation strategies for the Flan-T5 and Zephyr-7B models on the SST2 dataset. The ‘Random‘ strategy repeats the training and evaluation multiple times without any constraints. In the ‘Fixed‘ strategy, the randomness factor configuration is kept fixed to a single state during the investigation. Our method (‘Interactions‘) takes the interactions into consideration. Factors considered important for different strategies are denoted using the (*) symbol. We observe that interactions between factors may cause some factors to have their importance overestimated (denoted in bold) or underestimated (denoted in italics).
4.1 Interactions Between Randomness Factors
In this section, our goal is to answer the following research question: RQ1: How do the interactions between randomness factors affect their individual importance? To answer this question, we compare our proposed method (Interactions) with the commonly used investigation methods: 1) Random, which varies the overall random seed in the investigation without any constraint on the configurations of other factors; and 2) Fixed, where the non-investigated randomness factors are fixed to a single configuration for all runs of the investigation. For these strategies, we consider the effects of factor to be important when it contributes at least 50% of the golden model standard deviation. The results from this comparison are shown in Table 1, with detailed results in Appendix E.
Effects of randomness factors may be overestimated or underestimated when interactions are not taken into consideration. The ‘Random‘ strategy leads to a deviation similar to the one from the golden model across all investigated randomness factors. Such result indicates that all randomness factors are equally important, leading to a significant importance overestimation in some cases (e.g., Data Order factor for both Flan-T5 and Zephyr models). Even though the ‘Fixed‘ strategy produces more reliable results, it is still affected by the random choice of the single factor configuration (e.g., Data Order contributing deviation of $3.014$, which is much higher than the deviation of $2.244$ from the golden model). As such, we observe underestimation of the results (e.g., Sample Choice and Data Split with the Zephyr-7B model not being considered important with a deviation of $0.406$ and $0.402$) as well as overestimation (e.g., Data Order being considered important with a deviation of $3.014$ for Flan-T5 and $0.957$ for Zephyr-7B). Taking the interactions into consideration, we observe that the Data Order randomness factor is not consistently important for in-context learning even when choosing samples randomly, which confirms the impact of interactions on incorrect attribution of effects of different randomness factors.
4.2 Importance of Randomness Factors
4.3 Effects of Variable Number of Classes and In-Context Samples
In this section, our focus is on answering the following research question: RQ3: How does the importance of data-specific randomness factors change based on the number of classes and in-context samples? As we observe different effects of randomness factors on binary and multi-class datasets, our main focus is to determine whether the change in importance is caused by the increased number of in-context examples in the prompt, by the larger number of options that can be predicted or by a combination of both. The results from changing the number of classes are included in Figure 1, and the results from changing the number of shots for in-context learning are included in Figure 2.
The importance of the Sample Choice for in-context learning is not consistently affected by the number of classes. In case of Flan-T5 and Zephyr, the importance of Sample Choice gradually decreases as we increase the number of classes (from $0.57$ on SST2 to $-0.57$ on SNIPS for Flan-T5, or from $0.39$ on SST2 to $0.17$ on DB-Pedia for Zephyr). For the LLaMA-2 model, the decrease is not as consistent, with the importance being much lower on the TREC dataset than on the SNIPS dataset. Finally, the Sample Choice randomness factor is consistently important across all datasets for the Mistral model, with no apparent tendency.
4.4 Impact of Prompt Format
5 Conclusion
In this work, we have proposed a novel method that explicitly takes interactions between different randomness factors into consideration by mitigating the effects of the other, non-investigated, randomness factors. In addition, our method is designed to determine the importance of the investigated randomness factor by measuring what fraction of the overall deviation of the model (represented using a golden model) it contributes over the mitigated randomness factors.
Applying our proposed method to investigate the effects of randomness factors on the stability of in-context learning, fine-tuning and meta-learning, we confirm our hypothesis that interactions between randomness factors may cause incorrect attribution of effects of one factor to another, leading to inconsistent result. Contrary to the previous works, after taking interactions into consideration, we do not observe a consistent sensitivity of in-context learning approaches to the sample order even when choosing samples at random. Instead, we observe that the importance of randomness factors, especially the sample order, is affected by the interactions with other effects of randomness, as well as by the more systematic choices such as the number of predicted classes, the number of samples per class and the choice of prompt format.
These findings highlight the need for more systematic investigation and design of new mitigation strategies to stabilise the approaches by reducing the effects of randomness (such as those designed in related work Zhang et al. (2022); Yang and Kwok (2022); Köksal et al. (2023); Li and Qiu (2023); Chang and Jia (2023)), especially when it comes to under-researched randomness factors and approaches for learning with limited labelled data.
Limitations
The effects of randomness factors are investigated on a selection of models from different approaches for learning with limited labelled data. However, in order to provide a more extensive and in-depth analysis of the interactions and the more systematic choices without a significant increase in computation costs, the effects are investigated on models of smaller sizes – we use the base versions of BERT, RoBERTa and Flan-T5 models, and a 4-bit quantised versions of the LLaMA-2-13B, Mistral-7B and Zephyr-7B models. As such, the observed effects may not be as representative for larger models. However, similar to related work, we observed the larger models to be more susceptible to the effects of randomness and so the results of our investigation may under-estimate the importance of different factors (instead of their over-estimation).
The number of investigation and mitigation runs used in our investigation is selected based on an ablation study (in Appendix 3). In addition, following related work (e.g., Gao et al. (2021); Chang and Jia (2023); Sclar et al. (2023); Li and Qiu (2023); Köksal et al. (2023)) and the results of our ablation study, we also evaluate each run using only 1 000 test samples. In both cases, the decision represents a trade-off between the reliability of the results and their feasibility. Increasing the number of runs and the number of test samples could potentially lead to more precise results (especially on larger datasets), but at the cost of significant computation costs increase. At the same time, the number of investigation and mitigation runs utilised in this paper still represents a significant improvement over the existing studies, as it is a common practice to investigate the effects using very low numbers of runs. As future work, we plan to explore the possibilities for effective mitigation strategies for all the randomness factors to mitigate their effects and further reduce computation costs.
Acknowledgements
This work was partially supported by the projects funded by the European Union under the EU Horizon 2020: TAILOR, GA No. 952215, by the European Union under the Horizon Europe: DisAI, GA No. 101079164 and vera.ai, GA No. 101070093.
Appendix A Ethical Considerations and Impact Statement
Impact Statement: CO2 Emissions Related to Experiments
[…]
Appendix B Formal Description and Additional Resources Describing the Proposed Investigation Method
@@
In this Appendix, we provide the formal description of the proposed method, along with additional resources that should allow for more easier understanding of how the method operates. The proposed investigation method is designed for investigating effects of any randomness factor, while explicitly taking interactions with effects of other factors into consideration, and measuring the importance of the found effects. Overall, the effects are investigated by observing how the performance changes across the different states the investigated randomness factors can appear in. To deal with the interactions, the effects of remaining randomness factors are mitigated (i.e., the deviation they contribute is reduced as much as possible). To determine the importance, we compare the contributed deviation of the investigated randomness factor with effects of other factors, and with the overall deviation from a golden model. The golden model represents the objective estimate of the performance deviation and is obtained by training a model while mitigating all the randomness factors at the same time. The final importance score is then determined as the fraction of the overall deviation (represented using the golden model) the investigated factor contributes over all the remaining, non-investigated factors.
The investigation steps are compiled in Algorithm 1, representing the pseudo code of the method as described in Section 3. The following section provides the formal description of the method (Appendix B.1), the illustration of how the method operates and the results it computes (Appendix B.2).
B.1 Formal Description of the Method
First, a set $\mathbb{RF}$ ($|\mathbb{RF}|=K$) is defined, which includes all the factors that will be considered in the investigation. Each randomness factor is characterized by a set of its randomness factor configurations, $\mathbb{C}_{j}$, specifying all the possible states the factor can appear in. For example, the different permutations of samples represent the configurations of the data order randomness factor.
To investigate randomness factor $i$, it is first separated from all the other randomness factors that will have their effects mitigated. An investigated factor configurations set $\mathbb{IFC}_{i}$ (the size of this set, $|\mathbb{IFC}_{i}|=N$, dictates the number of repeated runs in a single investigation run of the method) is defined as $\mathbb{IFC}_{i}=\mathbb{C}_{i}$, containing configurations used to observe the behaviour of the model. The proposed method is easily extensible to investigate effects of multiple randomness factors at the same time by simply changing the definition of $\mathbb{IFC}_{i}$. The configurations for the mitigated randomness factors are jointly used to define the mitigated factor configurations set $\mathbb{MFC}_{i}$ (the size of this set, $|\mathbb{MFC}_{i}|=M$, dictates the number of mitigation runs of the method), as a cross product between the sets of randomness factor configurations leaving out the investigated randomness factor ($\mathbb{C}_{i}$):

### v2

@@ Abstract
While learning with limited labelled data can effectively deal with a lack of labels, it is also sensitive to the effects of uncontrolled randomness introduced by so-called randomness factors (i.e., non-deterministic decisions such as choice or order of samples). We propose and formalise a method to systematically investigate the effects of individual randomness factors while taking the interactions (dependence) between them into consideration. To this end, our method mitigates the effects of other factors while observing how the performance varies across multiple runs. Applying our method to multiple randomness factors across in-context learning and fine-tuning approaches on 7 representative text classification tasks and meta-learning on 3 tasks, we show that: 1) disregarding interactions between randomness factors in existing works led to inconsistent findings due to incorrect attribution of the effects of randomness factors, such as disproving the consistent sensitivity of in-context learning to sample order even with random sample selection; and 2) besides mutual interactions, the effects of randomness factors, especially sample order, are also dependent on more systematic choices unexplored in existing works, such as number of classes, samples per class or choice of prompt format.
@@
Branislav Pecher, Ivan Srba, Maria Bielikova Faculty of Information Technology, Brno University of Technology, Brno, Czechia Kempelen Institute of Intelligent Technologies, Bratislava, Slovakia Slovak.AI, Bratislava, Slovakia {branislav.pecher, ivan.srba, maria.bielikova}@kinit.sk
@@ 1 Introduction
Learning with limited labelled data, such as in-context learning, fine-tuning or meta-learning, is an umbrella term for approaches designed to work when enough labels are lacking. Although such approaches can effectively deal with limited labels, they were observed to be notably sensitive to the effects of uncontrolled randomness. Such randomness is introduced by the randomness factors, which represent the non-deterministic decisions in the training process, such as order of samples, the model initialisation or sample choice Pham et al. (2021); Gundersen et al. (2022); Pecher et al. (2024b).
@@
The randomness in the training process can have massive impact, leading to large deviation in the performance over multiple training runs. In-context learning was found to be sensitive to the order of samples, where changing only order of samples leads from state-of-the-art predictions to random guessing Lu et al. (2022); Zhao et al. (2021b). Similarly, repeating fine-tuning and evaluation multiple times with different random seeds can result in smaller language models outperforming their larger counterparts Dodge et al. (2020). If the randomness is not properly addressed, it can have non-negligible negative consequences even with enough labelled samples Reimers and Gurevych (2017); McCoy et al. (2020). It was identified as a major obstacle to reproducibility Albertoni et al. (2023) that can prohibit objective comparison and cause a method to be incorrectly denoted as state-of-the-art only based on more favourable chance Reimers and Gurevych (2017). The uncontrolled randomness can also unintentionally (but, unfortunately, also intentionally by cherry-picking) create an imaginary perception of research progress.
@@
A lot of focus is dedicated to investigating and mitigating the effects of randomness and sensitivity of learning with limited labelled data Mosbach et al. (2021); Pecher et al. (2024b), especially for in-context learning Lu et al. (2022); Zhao et al. (2021b); Chang and Jia (2023); Li and Qiu (2023); Köksal et al. (2023). However, the existing research is often limited in its extent (in terms of randomness factors, approaches or settings) and at times leads to contradictory or inconsistent results. For example, in-context learning was believed to be consistently sensitive to the order of the randomly selected samples Lu et al. (2022); Zhao et al. (2021b), however, it was later observed that this sensitivity disappears when a more sophisticated sample selection strategy is used Zhang et al. (2022); Chang and Jia (2023); Li and Qiu (2023).
@@
We argue that the observed inconsistencies are caused by disregarding the interactions between randomness factors, which leads to incorrectly attributing the performance deviations to different randomness factors. Such interactions are so far partially or even completely overlooked in the existing works, resulting in the misleading findings. In addition, we hypothesise that the sensitivity of in-context learning is not only affected by the interactions with other factors but also by other systematic choices, which are not thoroughly controlled and explored in the existing works, such as the number of classes, shots per class and prompt format.
@@
Our main contributions are as follows [^1] To support replicability and extension of our results, we openly publish the source code of our proposed investigation method and experiments for determining the factor importance at https://github.com/kinit-sk/L3D-sensitivity-investigation:
@@
  • We propose a novel method for investigation of randomness factors’ effects that, in contrast to the existing works, is thoroughly formalised and explicitly addresses interactions between them by mitigating the effects of other non-investigated factors. In addition, it measures the relative importance of factors, by calculating what fraction of the overall deviation in the performance (estimated by a golden model) the investigated factor contributes in comparison to all other factors, which allows for more in-depth analysis across factors, models, datasets and experimental settings.
@@
  • We further analyse how the more systematic choices influence the importance of the randomness factors. We find the following key insights: 1) predicting a higher number of classes leads to increased importance of sample order for in-context learning and reduced importance of sample order and model initialisation for fine-tuning approaches; 2) increasing the number of in-context samples reduces the importance of sample selection while having no consistent effect on the importance of sample order; and 3) the choice of prompt format has a significant impact on the importance of different factors, with larger models showing lower sensitivity to this choice.
@@ 2 Related Work
The main strategy for investigating the effects of randomness factors is to repeat the training and evaluation multiple times, changing specific non-deterministic decisions of the training, such as changing what data is used and observing the change in results (i.e., Random strategyMcCoy et al. (2020); Dodge et al. (2020); Bouthillier et al. (2019); Agarwal et al. (2021). As such investigation may be affected by the interactions with other factors, another possibility is to perform the investigation by fixing all the other factors to a specific state (i.e., Fixed strategy), either chosen randomly Boquet et al. (2019); Pham et al. (2021); Zhao et al. (2021a) or as a result of a mitigation strategy Li and Qiu (2023); Chang and Jia (2023). Another investigation strategy is to vary all the investigated factors at the same time and then decouple their effects in evaluation Dodge et al. (2020); Bouthillier et al. (2021); Sellam et al. (2022); Weber et al. (2023); Webson and Pavlick (2022), which accounts for the interactions but introduces a significant increase in computation costs Bouthillier et al. (2021).
@@
Majority of the focus on investigating and mitigating effects of randomness is on in-context learning, which was found to be especially sensitive to the choice of samples Liu et al. (2022); Zhang et al. (2022); Chang and Jia (2023); Li and Qiu (2023); Köksal et al. (2023) and their order Lu et al. (2022); Zhao et al. (2021b); Nguyen and Wong (2023). However, it was observed that the sensitivity to sample order disappears when using a more sophisticated sample selection strategy instead of random selection Zhang et al. (2022); Chang and Jia (2023); Li and Qiu (2023), hinting at interactions between these factors that may lead to inconsistent results. In addition, the performance of in-context learning was found to be sensitive to more systematic choices as well Weber et al. (2023), such as the format of the prompt Sclar et al. (2023); Voronov et al. (2024) or number of shots Liu et al. (2022); Mavromatis et al. (2023). However, the impact of these systematic choices on the effects of randomness factors is not thoroughly investigated. Besides order of in-context examples, large language models were found to be especially sensitive to the order of choices in multi-choice question answering (Zong et al., 2023; Wei et al., 2024). Although the remaining approaches and randomness factors receive only limited focus, they were still found to be sensitive to the effects of randomness, such as fine-tuning being sensitive to the random seeds (that influence model initialisation and order of samples) Dodge et al. (2020); McCoy et al. (2020); Mosbach et al. (2021); Zhao et al. (2021a); Zhong et al. (2021), meta-learning being sensitive to the choice of adaptation samples or how they are split into tasks Agarwal et al. (2021); Setlur et al. (2021); Cioba et al. (2022); Ye et al. (2021), or the overall machine learning being sensitive to factors such as the impact of framework and hardware implementation Boquet et al. (2019); Pham et al. (2021), or the data split Bouthillier et al. (2019, 2021).
@@
In majority of the cases, the effects of randomness are evaluated based on a single aggregated metric from multiple runs (e.g., mean, standard deviation, or the difference between best and worst run), with the importance being determined in a binary fashion by comparing this metric to a threshold, which allows only for simple analysis McCoy et al. (2020); Ye et al. (2021); Zhang et al. (2022)). A slightly more nuanced analysis is possible only in specific cases, where statistical approaches are used, such as grouping runs and aggregating on group level Dodge et al. (2020), decoupling interactions Boquet et al. (2019) or estimating distribution from lower number of training runs Sellam et al. (2022). However, almost no studies analyse the importance of the effects in a way that would allow for easy comparison across different settings, such as what fraction of the overall variance the specific factor contributes.
@@ 3 Investigation of Randomness while Taking Interactions into Consideration
We propose a new method for investigating the effects of any randomness factor that takes the interactions between the effects of other factors into consideration, and which is designed to measure the importance of the found effects. The steps of the method are compiled in Algorithm 1, with further supplementary details included in Appendix B).
Setup.
First, a set $\mathbb{RF}$ ($|\mathbb{RF}|=K$) is defined, which includes all the factors that will be considered in the investigation. Each randomness factor is characterised by a set of its randomness factor configurations, $\mathbb{C}_{j}$, specifying all the possible states the factor can appear in. For example, the different permutations of samples represent the configurations of the data order randomness factor. For each factor $i$, the investigated factor configurations set $\mathbb{IFC}_{i}$, containing the configurations used for the investigation, is defined as $\mathbb{IFC}_{i}=\mathbb{C}_{i}$, and the mitigated factor configurations set $\mathbb{MFC}_{i}$, containing the joint configurations of the remaining randomness factors, is defined as a cross product between all the sets of randomness factor configurations, except for the investigated randomness factor ($\mathbb{C}_{i}$):